OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ 12y ^{2} }{ -18y }\]

rvc (rvc):

see if you can divide further

OpenStudy (anonymous):

i got \[\frac{ 2y }{ -3 }\]

rvc (rvc):

correct :)

OpenStudy (anonymous):

Not according to the answer key. \[ \frac{ -2y }{ 3 }\]

rvc (rvc):

we can write the negative sign as :\[\huge - \frac{2y}{3}\]

OpenStudy (anonymous):

Since the answer key has -ve sign next to 2, wouldn't it be wrong if we place it oustide? As in, wouldn't it be different ?

rvc (rvc):

it wont make any difference :)

OpenStudy (anonymous):

Ok, Thank you

rvc (rvc):

Thank you so much :)